Question 5: Below is the chemical equation catalyzed by isocitrate dehydrogenase (i... the third reaction in the citric acid cycle). α-ketoglutarate + CO2 + NADH + H+ isocitrate NAD (a) Calculate the physiological ΔG of the isocitrate dehydrogenase reaction at 25°C and pl 17.0 given: [NAD+]IN ADH]-8; [α-ketoglutarate] = 0.1 mM: [isocitrate] = 0.02 mM; standard state conditions for CO2 and H: and ΔG (b) Is this reaction a likely site for metabolic control? Explain. 。,--21.0 kJ/mol?
Answer #1

For the given reaction

a) dG = dGo + RT lnK

pH = -log[H+] = 7

[H+] = 1 x 10^-7 M

K = 0.1 x 1 x 10^-7 x 1/0.02 x 8 = 6.25 x 10^-8

dG = -21 x 1000 + 8.314 x 298 ln(6.25 x 10^-8) = -62.10 kJ/mol